# Neural Honeytrace: A Robust Plug-and-Play Watermarking Framework against Model Extraction Attacks

### v3

@@ Abstract
Developing high-performance deep learning models is resource-intensive, leading model owners to utilize Machine Learning as a Service (MLaaS) platforms instead of publicly releasing their models. However, malicious users may exploit query interfaces to execute model extraction attacks, reconstructing the target model’s functionality locally. While prior research has investigated triggerable watermarking techniques for asserting ownership, existing methods face significant challenges: (1) most approaches require additional training, resulting in high overhead and limited flexibility, and (2) they often fail to account for advanced attackers, leaving them vulnerable to adaptive attacks.
In this paper, we propose Neural Honeytrace, a robust plug-and-play watermarking framework against model extraction attacks. We first formulate a watermark transmission model from an information-theoretic perspective, providing an interpretable account of the principles and limitations of existing triggerable watermarking. Guided by the model, we further introduce: (1) a similarity-based training-free watermarking method for plug-and-play and flexible watermarking, and (2) a distribution-based multi-step watermark information transmission strategy for robust watermarking. Comprehensive experiments on four datasets demonstrate that Neural Honeytrace outperforms previous methods in efficiency and resisting adaptive attacks. Neural Honeytrace reduces the average number of queries required for a worst-case t-Test-based copyright claim from to with zero training cost. The code is available at https://github.com/NeurHT/NeurHT.
@@ 1 Introduction
With the growing scale of training dataset and model parameter, the development of high-performance deep learning models has become increasingly expensive. As a result, instead of releasing their models publicly, model owners usually build up Machine Learning as a Service (MLaaS) platforms for paid intelligent services, where users interact with black-box models via a query interface.
@@
However, malicious users can still extract valuable information by executing carefully-designed queries to the service interface, enabling them to locally reconstruct the functionality of the victim model with minimal overhead [1, 2, 3, 4, 5, 6], namely model extraction attacks (MEAs). In recent years, model extraction attacks have been extensively studied across various information channels [1, 2], leveraging different data sources [3, 6], and employing diverse learning strategies [7, 8]. Recent research has also investigated adaptive attacks targeting potential defenses [9, 10, 11]. These mechanisms use effective strategies to remove misleading perturbations or watermarks embedded in the output predictions, causing severe challenges to the copyright protection of MLaaS applications.
@@
To mitigate the risk of model extraction attacks, a variety of defense mechanisms have been proposed, including model extraction detection [5, 12], prediction perturbation [13, 14, 15, 11], and model watermarking [16, 17, 18, 19, 20]. Compared to the other two passive defenses, model watermarking aims to implant triggerable watermarks into the stolen model, allowing the model owner to assert ownership by activating the watermarks during inference. Recently, backdoor-like watermarks [18, 16, 20] have showed great potential for watermark embedding and capability retention. Fig. 1 provides an overview of the workflow of MEAs and triggerable watermarking methods, where the watermark is embedded in the output predictions and then transmitted to the stolen model.
Despite the success of previous solutions, existing watermarking strategies still face challenges in terms of flexibility and robustness. For example, the watermark embedding process in existing methods [18, 16, 20] requires extensive model retraining. Once embedded, these watermarks cannot be easily removed or modified. Some methods, including the state-of-the-art one, MEA-Defender [20], are designed for soft-label black-box scenarios, resulting in significant performance degradation in hard-label settings—the most common scenario in MLaaS. Meanwhile, existing approaches are only evaluated against naive attacks, where attackers have no knowledge about potential defenses. As shown in Sec. 5.2, these methods exhibit limited robustness against adaptive MEAs. More importantly, the rationale for why existing methods work and fail in different scenarios remains unclear, making it difficult to design flexible and robust watermarking methods. Tab. I compares the capability of different watermarking strategies.
@@
Therefore, we propose Neural Honeytrace, a robust plug-and-play watermarking framework against model extraction attacks. We begin by developing a watermark transmission model from an information-theoretic perspective. Based on this framework, we draw three key conclusions: (1) the rationale for triggerable watermarks lies in the transmission of similarity between input samples and watermarks, (2) existing methods fail in hard-label scenarios due to the channel capacity being lower than the source rate, and (3) adaptive attacks add noise to the watermark transmission process and compromise the watermark when channel noise exceeds the tolerance limit. Building on these insights, we introduced: (1) a training-free watermarking method for plug-and-play and flexible watermarking by directly embedding similarity scores into output logits, and (2) a multi-step watermark information transmission strategy for robust watermarking by strategically label-flipping. The main contributions of this paper are summarized as follows:
@@
We propose Neural Honeytrace, the first training-free triggerable watermarking framework. It is designed as a plug-and-play manner, offering the flexibility for seamless removal or modification post-deployment.
@@
We establish a watermark transmission model using the information theory, which addresses several open questions pertaining to triggerable watermarking, including the nature of watermarking information and the factors affecting the success rate of watermarking transmission. Guided by the model, we introduce two watermarking strategies which enable training-free watermark embedding and robust transmission.
@@
We evaluate Neural Honeytrace against various model extraction attacks, including adaptive attacks with advanced attackers, and empirically show that Neural Honeytrace achieves significant better robustness and less overhead compared to existing methods. Compared with existing methods, Neural Honeytrace reduces the average number of samples required for a worst-case t-Test-based copyright claim from to with zero training cost.
@@ 2.1 Model Extraction Attack
The goal of model extraction attacks is to rebuild the functionality of the victim model locally. To achieve this object, attackers first collect or synthesize an unlabeled surrogate dataset , and then utilize the victim model to label the surrogate dataset and get the corresponding label set . The type of label is determined by the setting of MLaaS (e.g., scores, probabilities, and hard-labels). Subsequently, attackers can train a surrogate model following the optimization problem:
where is the loss function for evaluating the distance between the outputs of and . After training, the surrogate model should have similar functionality with the victim model, thus attackers can provide stolen services as shown in Fig. 1.
The basic model extraction attack defined above may fail under perturbation-based defenses [13, 14, 15, 11]. Therefore, attackers introduced more advanced adaptive attacks to bypass potential defenses. Denote the perturbed prediction set as , where represents the perturbation added on by defense strategies. Adaptive model extraction attacks adopt different recovery mechanisms to recover from . For example, Smoothing Attack [9] performs several image augmentations on each query image and leverages the average of all predictions to recover the clean labels.
@@ 2.2 Triggerable Watermarking
Triggerable watermarking strategies [18, 16, 19, 20] aim at implanting watermarks that can survive during model extraction attacks into the target model. Similar to backdoor activation in backdoor attacks, model owners can use predefined triggers to activate watermarks in stolen models and utilize special model outputs for ownership declaration. Existing triggerable watermarking methods follow the injection of backdoors, for a predefined watermark trigger , the watermarking process can be described as the following optimization problem:
where is the watermarked model, are the sets of training data and labels, is the weight parameter which balances the watermarking success rate and clean accuracy, represents the watermark injection mechanism, and denotes the special output for IP declaration.
A fundamental challenge for triggerable watermarks is they need to survive during model extraction, where the surrogate dataset may not contain the watermark trigger predefined by defenders. Existing methods introduced different assumptions and addressed this challenge empirically by adding regularization terms to Eq. 2.2 [16, 19] or designing specialized triggers [18, 20].
@@ 2.3 Hypothesis Test
Although triggerable watermarking enables defenders to claim ownership by querying the stolen model, a single query is insufficient for the precision requirement (e.g., a false positive rate less than 0.0001). Therefore, it is necessary to conduct multiple experiments and analyze them in terms of statistical significance.
Previous methods [16, 20] adopt t-Test in Hypothesis Test to help achieve ownership statements. Specifically, a t-Test helps determine whether the observed difference in model behavior (Watermark Success Rate, WSR) is statistically significant, considering variability in the data. The hypothesis test can be formulated as:
Null Hypothesis: : The model does not contain the watermark (i.e., the difference between observed and expected success rate is due to chance).
Alternative Hypothesis: : The model contains the watermark (i.e., the difference is statistically significant and not due to chance).
To achieve a statistically significant result, a minimum sample size is required to ensure the power of the test is adequate to detect differences between the null and alternative hypotheses. The required sample size can be calculated using the following equation:
where is the critical value of the significance level, is the power, and (WSR) is the effect size. For example, assuming that the acceptable false positive rate and false negative rate are and , respectively, it will require queries to the target model to achieve copyright statement with WSR. As depicted in Fig. 2, the number of samples required grows exponentially as the success rate of watermarking decreases.
Intuitively, hypothesis testing implements trading query numbers for accuracy. As the result, the validity of watermarking depends on its ability to accomplish hypothesis testing with a limited number of queries given the WSR.
@@ 2.4 Threat Model
Attacker’s goal. As defined in Sec. 2.1, the primary goal of model extraction attackers is to reconstruct the functionality of the victim model locally. We assume the victim model is deployed as a black-box service, which means that attackers can only access output predictions through the interface provided by the model owner.
Attacker’s capability. We consider attackers who lack direct access to the training dataset but possess knowledge of the task domain (e.g., image classification, face recognition etc.) and utilize open-source datasets from the corresponding domains to perform attacks.
To account for varying levels of sophistication, we classify attackers into three categories based on their knowledge of potential defenses: 1) naive attackers who have no prior knowledge, 2) adaptive attackers who know the type of deployed defenses but do not know the implementation details, and 3) oracle attackers who have full knowledge about the deployed defense mechanisms.
Defender’s Capability. We assume that the defender does not have control over the model training process and provides defense services after the target model has been deployed. The defender possesses a watermark dataset along with the corresponding watermark features extracted by the target model. During the watermark embedding process, the defender can modify the predictions of the target model. For watermark activation, the defender interacts with the suspicious model with limited query numbers.
@@ 3 Watermark Transmission Model
Previous studies [16, 18, 19, 20] have empirically tackled challenges associated with triggerable watermarks. However, some fundamental questions remain unexplored:
Why are watermarks transmitted to stolen models even when they are not activated during MEAs?
Why do existing methods fail under hard-label scenarios or against adaptive attacks?
By examining the watermark transmission process, we identify its primary objective: to transfer specific information from the watermarked model to the stolen model. This process can be modeled as a message transmission problem. Accordingly, we develop a watermark transmission model grounded in information theory to answer the questions.
3.1 Elements
Initially, we modeled the watermark transmission process following the classic message transmission model. As illustrated in Fig. 3, the model owner designs a watermark represented by the message . This message is then encoded by an encoder, denoted as , according to a predefined strategy. The encoded message is transmitted via the model outputs, which serve as the communication channel. If the watermark message is successfully transmitted, it becomes embedded in the stolen model.
According to the model, three elements determine the performance of watermark transmission:
Source. The format of watermark message determines the amount of information contained in it, which can be quantified using information entropy:
where is the probability distribution of .
Encoding. In watermark transmission scenarios, the encoding process determines how watermarks change model outputs, which can described as:
where is the alphabet for encoding, and denote the modified and the original output, respectively.
Channel. According to Shannon’s theorem [21], the channel capacity can be calculated using mutual information :
For watermark transmission, and both represent the output probability, thus , and .
3.2 Theoretical Analysis
For triggerable watermarks [16, 18, 19, 20], the watermark message can not be ”Does the sample contain the trigger?” because triggers are not distributed in the surrogate dataset, making bit during MEAs. Instead, triggerable watermarking methods utilize the similarity scores to indirectly transmit watermark information.
Specifically, previous work [22] observes the long-tailed effect in backdoor learning, where clean samples with characteristics similar to backdoor samples are more likely to output the backdoor class. We observe that this effect can also be exploited in reverse for watermarking (backdoor) implantation, i.e., by adjusting the probability of the target class based on the similarity scores, making the target model inherit the watermarks.
In previous triggerable watermarking methods [16, 18, 19, 20], the long-tailed effect is constructed in the protected model during watermark implantation and transmitted to the stolen model during MEAs. As depicted in Fig. 4, by gradually replacing the corresponding region with the watermark trigger, the probability of the target class gradually increases (gray line). A similar trend can be observed on extraction query examples (blue plots), too. As the result, the stolen model will learn the relationship between trigger similarity and output probability.
Based on Q1, we provide an running example to show why existing methods succeed or fail in different scenarios: According to the Central Limit Theorem, the similarity distribution approximately follows a normal distribution, , the information entropy is given by , where is determined by the query data distribution and the similarity calculation method. For instance, in Fig. 4, setting the precision of similarity values to results in an estimated bit.
In soft-label black-box scenarios, we can estimate the channel capacity using Eq. 3. For example, in Fig. 4, if the precision of output probabilities is set to , the channel capacity is approximately . By Shannon’s theorem [21], given an acceptable error rate , there exists an encoding strategy that ensures:
where and represents the maximum transmission rate. We observe that if , there exists a encoding strategies that makes , which means the watermark information can be transmitted correctly with close to 100% probability. Furthermore, for a balanced dataset, the label information entropy is approximately . In Fig. 4, and since , watermark transmission will not compromise the transmission of label information, i.e., model usability.
However, for hard-label scenarios, the maximum channel capacity satisfies . In Fig. 4, , according to Eq. 4, the error rate , which means the channel is not able to transmit watermark information with a high probability. Moreover, even if the value of becomes large enough for transmitting watermark information, the channel is exclusive, i.e., only one of the watermark or label information can be transmitted through a single query. In this case, the stolen model can not receive enough information for watermark implantation.
For adaptive MEAs [7, 9, 10, 11], without loss of generality, the watermark removal mechanisms can be observed as adding channel noises in watermark transmission. In this case, the channel capacity can be calculated as:
where B is the channel bandwidth, and
where are the mean and variance of the signal and noise, respectively. For MEAs, the mean of the noise added by attackers is approximately equal to because they have no prior knowledge about the label information, and a mean greater than will introduce bias and leads to failure of the attack. Therefore, the channel capacity and the variance of noise are negatively correlated in a logarithmic relationship. Once the noise variance increases to the point where the corresponding channel capacity is less than the source rate, the adaptive MEAs successfully invalidate watermark transmission.
4 Neural Honeytrace Design
In this section, we introduce our watermarking framework, Neural Honeytrace. We first provide the brief workflow of Neural Honeytrace in Sec. 4.1. Then we provide detailed description of training-free watermark embedding and multi-step watermark transmission in Sec. 4.2 and Sec. 4.3.
4.1 Overview
Based on the theoretical analysis, we design a robust plug-and-play watermarking framework, Neural Honeytrace. Specifically, Neural Honeytrace addresses the following challenges faced by existing watermarking methods:
Flexibility Based on Q1-A1, the transmitted watermark information is the similarity score, thus we design Neural Honeytrace as a plug-and-play manner by directly calculating the feature similarity score and embedding it into output logits, offering the flexibility for seamless removal or modification post-deployment.
Robustness Based on Q2-A2 and Q3-A3, limited channel capacity and noise tolerance lead to transmission failure, thus we introduce the similarity-guided label flipping strategy, which embeds watermark information in label distributions and enables multi-step transmission, improving transmission success rates.
Fig. 5 provides an overview of the workflow of generic model extraction attacks and Neural Honeytrace. For Neural Honeytrace, the watermarking process consists of four steps. Step 1: defenders initially select watermarks from the watermark pool (e.g., random watermark, semantic watermark, and composite watermark etc.). Step 2: the target model extracts the watermark features and query features respectively and makes predictions on the query samples. Step 3: Neural Honeytrace calculates the similarity of these two features and embeds the similarity into the predictions by mixing the reference logits (target logits of watermarking) and the original logits following the similarity value. Step 4: Neural Honeytrace establishes a probability-guided label-flipping matrix according to the similarity. By encoding similarity values into the distribution of different queries, this matrix enables multi-step watermarking transmission, which is robust under hard-label scenarios and against adaptive attacks.
For ownership statement, defenders use watermarked samples to query the suspicious model and perform a hypothesis test. According to Eq. 2, given the watermark success rate, if the total number of watermarked samples is lager than the lower bound, then the suspicious model can be considered statistically significant as having a watermark.
@@ 4.2 Training-free Watermark Embedding
As we analyzed in Sec. 3.2, the information transmitted by triggerable watermarks is the similarity of query features and watermark features, and the message channel is the output predictions. Therefore, we can directly calculate and encode the similarity into predictions without additional training. So the initial questions become watermark selection, similarity calculation, and similarity embedding.
Watermark generation. The form of watermarks will influence the cost of watermarking process. Previous methods have introduced different watermark generation strategies, e.g., EWE [16] utilized white pixel blocks as triggerable watermarks, while MEA-Defender [20] used spliced in-distribution samples as triggers.
Intuitively, label-independent information will gradually lost during the forward process of neural networks (information bottelneck [23]). And if the watermark only contains out-of-distribution features (e.g., white pixel blocks), it will introduce additional information for the stolen model to learn and make the watermark transmission process harder to converge. Therefore, Neural Honeytrace adopts composite in-distribution samples as watermarks following MEA-Defender [20]. We also evaluated Neural Honeytrace with different watermark forms in Sec. 5.3.
@@
where is a hyperparameter which balances the watermark success rate and model usability, represents the last layer of the target model, and denotes the i-th watermark. The calculated distance is clipped to using CLIP.
Additionally, considering that the attackers probably will not have a large amount of in-distribution data for querying the target model, we adopt a simple algorithm to minimize the impact of watermarking on model availability as follows:
Similarity embedding. Given the similarity, Neural Honeytrace uses logits-mixing to embed the similarity information into the predictions. Denotes the original logits as , the reference logits of the target watermark class as , then the mixed logits can be calculated as:

### v4

# Neural Honeytrace: Plug&Play Watermarking Framework against Model Extraction Attacks

@@ Abstract
Triggerable watermarking enables model owners to assert ownership against model extraction attacks. However, most existing approaches require additional training, which limits post-deployment flexibility, and the lack of clear theoretical foundations makes them vulnerable to adaptive attacks. In this paper, we propose Neural Honeytrace, a plug-and-play watermarking framework that operates without retraining. We redefine the watermark transmission mechanism from an information perspective, designing a training-free multi-step transmission strategy that leverages the long-tailed effect of backdoor learning to achieve efficient and robust watermark embedding. Extensive experiments demonstrate that Neural Honeytrace reduces the average number of queries required for a worst-case t-test-based ownership verification to as low as of existing methods, while incurring zero training cost.
@@ 1 Introduction
Malicious users can locally reconstruct the functionality of the victim model by executing carefully-designed queries to the Machine Learning as a Service (MLaaS) interface, namely model extraction attacks (MEAs) (Tramèr et al., 2016; Orekondy et al., 2019; Juuti et al., 2019; Truong et al., 2021). In recent years, model extraction attacks have been extensively studied across various information channels (Tramèr et al., 2016), leveraging different data sources (Truong et al., 2021), and employing diverse learning strategies (Jagielski et al., 2020; Chandrasekaran et al., 2020). Recent research has also investigated adaptive attacks targeting potential defenses (Lukas et al., 2022; Chen et al., 2023; Tang et al., 2024).
@@
To mitigate the risk of model extraction attacks, a variety of defense mechanisms have been proposed, including model extraction detection (Juuti et al., 2019; Kesarwani et al., 2018), prediction perturbation (Kariyappa and Qureshi, 2020; Lee et al., 2019; Mazeika et al., 2022; Tang et al., 2024), and model watermarking (Jia et al., 2021; Szyller et al., 2021; Lin et al., 2020; Cong et al., 2022; Lv et al., 2024b). Compared to the other two passive defenses, model watermarking aims to implant triggerable watermarks into the stolen model, allowing the model owner to assert ownership by activating the watermarks during inference. Recently, backdoor-like watermarks (Lin et al., 2020; Jia et al., 2021; Lv et al., 2024b) have showed great potential for watermark embedding and capability retention.
Despite the success of previous solutions, existing watermarking strategies still face challenges in terms of flexibility and robustness. For example, the watermark embedding process in existing methods (Lin et al., 2020; Jia et al., 2021; Lv et al., 2024b) requires extensive model retraining. Once embedded, these watermarks cannot be easily removed or modified. Meanwhile, existing approaches are only evaluated against naive attacks and suffer from unacceptable query overhead when faced with adaptive attacks. As shown in Tab. 1, model owners need to query the suspicious model many times for asserting ownership.
@@
Therefore, we propose Neural Honeytrace, a plug-and-play watermarking framework against model extraction attacks. We begin by developing a watermark transmission model from an information-theoretic perspective. Based on this framework, we draw two key conclusions: (1) the rationale for triggerable watermarks lies in the transmission of similarity between input samples and watermarks, (2) existing methods fail in adaptive attack scenarios due to the channel capacity being lower than the source rate. Building on these insights, we introduced a training-free multi-step method for robust and flexible watermarking. The main contributions of this paper are summarized as follows:
@@
We propose Neural Honeytrace, a training-free triggerable watermarking framework, offering the flexibility for seamless removal or modification post-deployment.
@@
We establish a watermark transmission model using the information theory. Guided by the model, we introduce two watermarking strategies which enable training-free multi-step watermark transmission.
@@
Neural Honeytrace reduces the average number of samples required for a worst-case t-Test-based ownership assertion to as low as of existing methods with zero training cost.
@@ 2.1 Model Extraction Attack
Model extraction attacks aim to locally replicate the functionality of a victim model . Attackers first collect or synthesize an unlabeled surrogate dataset , query to obtain labels (e.g., scores, probabilities, or hard labels), and then train a surrogate model by minimizing
where measures the discrepancy between the outputs of and .
Perturbation-based defenses (Kariyappa and Qureshi, 2020; Lee et al., 2019; Mazeika et al., 2022; Tang et al., 2024) corrupt model outputs to disrupt this process. Let denote the perturbed predictions, where is the defense-induced perturbation. To bypass such defenses, adaptive attacks introduce recovery mechanisms to infer clean labels from . For instance, the Smoothing Attack (Lukas et al., 2022) averages predictions over multiple input augmentations to suppress output perturbations.
@@ 2.2 Triggerable Watermarking
Triggerable watermarking (Lin et al., 2020; Jia et al., 2021; Cong et al., 2022; Lv et al., 2024b) embeds verifiable watermarks into a model such that they can be activated in extracted copies. Similar to backdoor attacks, predefined triggers are used to elicit distinctive outputs for ownership verification. Most existing methods inject watermarks during training by optimizing
where denotes the watermarked model, applies the trigger , and is the watermark-specific output.
@@ 2.3 Hypothesis Test
Since a single trigger query cannot satisfy stringent precision requirements (e.g., false positive rate ), ownership verification relies on statistical hypothesis testing over multiple queries. Prior works (Jia et al., 2021; Lv et al., 2024b) employ a t-test on the watermark success rate (WSR) to assess whether the observed behavior is statistically significant. Specifically, the null hypothesis assumes the absence of a watermark, while the alternative hypothesis indicates its presence. To ensure sufficient test power, a minimum number of queries is required, given by
where and denote the false positive and false negative rates, respectively, and (WSR) is the effect size. For example, with and , achieving WSR requires queries. As shown in Fig. 1, the required sample size grows rapidly as WSR decreases.
@@ 2.4 Threat Model
Attacker The attacker aims to reconstruct the functionality of a victim model deployed as a black-box service, with access limited to model outputs. While lacking the victim’s training data, the attacker knows the task domain and leverages publicly available datasets for model extraction. We consider three attacker types: naive (no defense knowledge), adaptive (aware of the defense type), and oracle (full defense knowledge).
Defender The defender operates post-deployment and cannot modify the original training process. The defender maintains a watermark dataset and corresponding watermark features, and can alter model predictions to embed watermarks. Ownership verification is performed by querying a suspicious model under a limited query budget.
@@ 3 Watermark Transmission Model
Despite the initial success of existing triggerable watermarks (Jia et al., 2021; Lv et al., 2024b), two critical questions remain unanswered:
How are these watermarks transferred to the student model without being activated?
Why do they fail in certain scenarios (e.g., label-only scenarios or adaptive attacks (Lukas et al., 2022))?
Insight 1: Triggerable watermark transmission is a reverse application of the long-tailed effect.
Prior work on backdoors observes a long-tailed effect: clean samples that are semantically similar to backdoor samples exhibit elevated target-class confidence (Xu et al., 2024).
During model extraction, watermark triggers are absent from surrogate data, making direct transmission infeasible. We show that triggerable watermarking leverages the long-tailed effect in reverse. By embedding a smooth relationship between trigger similarity and output probability, the watermark becomes transferable without activation.
As shown in Fig. 2, the protected model with MEA-Defender watermarking exhibits a pronounced long-tailed effect on the test samples. Consequently, watermark transmission is achieved through probabilistic similarity embedding rather than discrete trigger activation.
Based on Insight 1, we model watermark implantation as a message transmission process as depicted by Fig. 3. A watermark message is implicitly encoded into model outputs via an encoding function , and the model output serves as the communication channel.
According to information theory (Shannon, 1948), the success rate of message transmission is determined by the source rate, coding scheme, channel capacity, and channel noise. Therefore, we arrive at Insight 2.
Insight 2: Triggerable watermarks fail when the source rate exceeds channel capacity.
We model watermark inheritance in black-box model extraction as a communication process with source rate , channel capacity , and transmission error . By Shannon’s theorem, reliable transmission is possible only if
where . In watermark transmission, random variable represents the output logits/label, and represents the received random variable after watermark embedding and potential label-recovery processes (i.e., ). According to the data processing inequality (Beaudry and Renner, 2012), we have
where is the information entropy of . In hard-label settings, task labels must always be transmitted. Thus, the source rate decomposes as where denotes watermark information (e.g., similarity-dependent outputs). However, according to Eq. 3, the channel capacity is fully consumed by label transmission (). The additional watermark rate therefore yields , leading to inevitable errors.
4 Neural Honeytrace
Based on Insights 1 and 2, we propose Neural Honeytrace, a robust plug-and-play watermarking framework. Fig. 4 illustrates the overall pipeline of Neural Honeytrace, which performs watermarking in four steps. (1) The defender samples watermarks from a predefined watermark pool. (2) Given a query, the target model extracts both watermark features and query features and produces the original prediction. (3) Neural Honeytrace computes the similarity between watermark and query features, and embeds this similarity into the output by interpolating reference logits with the original logits. (4) Based on the similarity, a probability-guided label-flipping matrix is constructed, which encodes similarity information into the output distributions across multiple queries, enabling multi-step watermark transmission.
For ownership verification, the defender queries a suspicious model using watermarked samples and conducts a hypothesis test. As shown in Eq. 2, if the number of queries exceeds the required lower bound under a given watermark success rate, the presence of the watermark can be established with statistical significance.
@@ 4.1 Training-free Watermark Embedding
As we analyzed, the information transmitted by triggerable watermarks is the similarity of query features and watermark features, and the message channel is the output predictions. Therefore, we can directly calculate and encode the similarity into predictions without training.
Watermark generation. The form of watermarks will influence the cost of watermarking process. Previous methods have introduced different watermark generation strategies, e.g., EWE (Jia et al., 2021) utilized white pixel blocks as triggerable watermarks, while MEA-Defender (Lv et al., 2024b) used spliced in-distribution samples as triggers.
Intuitively, label-independent information will gradually lost during the forward process of neural networks (information bottelneck (Tishby and Zaslavsky, 2015)). Therefore, Neural Honeytrace adopts composite in-distribution samples as watermarks following MEA-Defender (Lv et al., 2024b).
@@
where is a hyperparameter which balances the watermark success rate and model usability, represents the last feature layer of the target model, and denotes the i-th watermark. The distance is clipped to using CLIP.
Additionally, considering that the attackers probably will not have a large amount of member-data for querying the target model, we adopt a simple algorithm to minimize the impact of watermarking on model availability as follows:
Similarity embedding. Given the similarity, Neural Honeytrace uses logits-mixing to embed the similarity information into the predictions. Denote the original logits as , the reference logits of the target watermark class as , then the mixed logits can be calculated as: